# Fine-tuning Diffusion Policies with Backpropagation Through Diffusion Timesteps

### v2

@@ 6 Experiments
In this section, we provide a comprehensive evaluation of NCDPO across a variety of challenging environments. We begin by detailing the experimental setup in Sec. 6.1, followed by results on continuous robot control tasks in Sec. 6.2 and discrete multi-agent coordination tasks in Sec.6.3. Finally, we conduct ablation studies in Sec. 6.4 to assess the robustness and of NCDPO.
@@ 6.1 Environmental Setup
Environments: OpenAI Gym locomotion.
Our first set of experiments involves testing NCDPO on a series of well-established locomotion benchmarks from OpenAI Gym [3], namely: Hopper-v2, Walker2D-v2, and HalfCheetah-v2. The pre-trained Diffusion Policies used in these experiments are trained from the D4RL "medium" dataset [6], which contains a diverse range of pre-recorded trajectories. For the fine-tuning process, we use dense reward.
Environments: Robomimic.
We further evaluate the performance of NCDPO on robotic manipulation tasks within Robomimic benchmarks [14]. The specific scenarios we consider include Lift, Can, Square, and Transport, varying in difficulty. To ensure temporal consistency in actions, we employ action chunking with size 4 for Lift, Can, and Square, and size 8 for Transport, following the setting in [20]. All tasks are fine-tuned using sparse rewards, which provide feedback only in the form of success or failure signals.
Environments: Google Research Football. To evaluate NCDPO’s capability in large discrete action spaces, we test on three Google Research Football scenarios requiring multi-agent coordination: 3 vs 1 with Keeper, Counterattack Hard, and Corner. Here we adopt a centralized control strategy, where actions for all agents are generated simultaneously using a single Diffusion Policy. The base Diffusion Policies are pre-trained to output one-hot vectors corresponding to ground-truth actions. To construct the pre-training dataset, we aggregate trajectories from multiple MLP-based policies with varying success rates.
@@ 6.2 Evaluation on Continuous Robot Control Tasks
We first evaluate NCDPO on continuous control tasks across two benchmarks: OpenAI Gym locomotion and Robomimic. In these environments, we compare NCDPO with DPPO [20]; DRWR and DAWR [20], based on reward-weighted regression [17] and advantage-weighted regression respectively [16]; DIPO [31], which employs action gradients as the score function for denoising steps; and Q-learning-based methods such as IDQL [8] and DQL [28].
From the experimental results shown in Fig. 3 and Fig. 4, we observe that NCDPO consistently achieves the strongest performance and exhibits robustness across all tasks. While DPPO, which is the best among the baseline methods, performs comparably to NCDPO in the Robomimic benchmark  [^1] Note that we tested performance using the latest version of DPPO (v0.8), which is about 2.5x sample efficient in Transport task as reported in the original paper., it lags behind in the OpenAI Gym locomotion environments. Other baselines generally underperform relative to DPPO. Notably, IDQL demonstrates strong performance on the first three Robomimic tasks but fails in the final one. In contrast, DQL suffers from instability across all scenarios. We additionally conducted experiments on Square using vision-based inputs, with results provided in the appendix. The results demonstrate consistent improvements in Diffusion Policy performance when fine-tuned with NCDPO.
Figure 3: Performance comparison on OpenAI Gym locomotion tasks. Results are averaged over three seeds. NCDPO (ours) achieves the strongest performance.
Figure 4: Performance comparison on Robomimic tasks. Results are averaged over three seeds. NCDPO (ours) achieves the strongest performance.
Scenario NCDPO (Ours) DPPO AWR IDQL DQL RWR DIPO
Hopper-Medium 126.4 (1.8) 98.4 (2.0) 44.8 (1.2) 113.8 (0.2) 122.7 (1.2) 100.9 (3.6) 94.4 (4.9)
Hopper-Medium-Replay 128.2 (2.8) 114.6 (3.3) 82.1 (7.0) 117.9 (1.1) 121.1 (3.6) 104.2 (3.1) 108.6 (1.0)
Hopper-Medium-Expert 135.2 (2.0) 102.4 (5.2) 40.7 (0.5) 131.9 (0.4) 120.9 (28.1) 113.6 (4.7) 127.5 (1.3)
Walker2d-Medium 118.4 (3.8) 101.2 (1.6) 93.5 (8.3) 110.7 (0.5) 94.9 (36.9) 90.2 (3.3) 99.8 (2.0)
Walker2d-Medium-Replay 126.6 (4.5) 105.1 (4.3) 75.8 (2.3) 121.9 (2.6) 107.2 (37.8) 69.2 (6.0) 88.6 (11.0)
Walker2d-Medium-Expert 141.0 (2.3) 137.5 (2.0) 124.2 (5.4) 135.5 (2.7) 67.3 (47.6) 106.8 (6.8) 138.8 (1.4)
HalfCheetah-Medium 122.0 (11.0) 82.3 (0.7) 65.5 (2.9) 79.3 (0.8) 77.1 (5.4) 47.9 (5.4) 73.9 (0.6)
HalfCheetah-Medium-Expert 139.7 (6.8) 80.6 (1.6) 64.4 (2.3) 77.8 (1.0) 72.1 (8.0) 38.4 (2.9) 71.4 (0.2)
HalfCheetah-Medium-Replay 121.0 (2.0) 72.3 (0.4) 60.5 (0.7) 74.3 (0.9) 73.0 (3.1) 30.7 (1.8) 58.1 (0.9)
Lift 100.0 (0.0) 99.7 (0.2) 93.3 (1.7) 99.2 (0.1) 99.8 (0.3) 97.5 (0.5) 97.3 (0.8)
Can 99.3 (1.2) 99.0 (1.0) 33.8 (3.2) 94.5 (3.1) 0.3 (0.6) 90.7 (0.8) 52.8 (5.1)
Square 87.3 (4.5) 87.0 (2.3) 40.3 (8.5) 80.0 (5.0) 0.0 (0.0) 74.8 (2.5) 25.3 (4.5)
Transport 96.7 (2.31) 91.3 (2.9) 11.2 (3.5) 0.5 (0.8) 0.0 (0.0) 0.0 (0.0) 0.2 (0.3)
Table 1: Mean and standard deviation of performance over continuous robot control scenarios. Each result is evaluated on three different seeds. NCDPO (ours) exhibits the strongest performance. Performance on OpenAI Gym locomotion tasks are normalized according to scores of MLP policies trained from scratch using PPO with 1M samples reported in Tianshou [29] . Original scores are listed in Table 3.
@@ 6.3 Evaluation on Discrete Multi-agent Coordination Tasks
[…]
As Fig. 5 demonstrates, NCDPO outperforms the MLP baseline across all three evaluated scenarios. This outcome not only highlights the superiority of Diffusion Policy in handling complex and diverse demonstration data over simple MLP policy, but also confirms the effectiveness of the DP+NCDPO during the fine-tuning phase.
@@
6.4 NCDPO is Robust to the Number of Denoising Steps
We further conduct an ablation study to investigate the impact of varying the number of denoising steps in the diffusion model. The experimental results shown in Fig. 7 indicate that NCDPO demonstrates strong robustness to the choice of denoising steps.
Figure 6: Ablation Study on Denoising Steps in OpenAI Gym locomotion tasks.
Figure 7: Ablation Study on Denoising Steps in Robomimic tasks.
We hypothesize that the robustness of NCDPO arises from the way gradients are propagated through time during the diffusion process. This gradient flow leads to more accurate gradient estimates.
7 Conclusion and Limitations
We present NCDPO, a novel approach for fine-tuning Diffusion Policies through Proximal Policy Optimization that exhibits strong performance across continuous and discrete control domains. Our key innovation lies in reformulating the diffusion denoising process as a noise-conditioned stochastic policy that enables effective gradient backpropagation through diffusion timesteps. Through extensive experiments across locomotion, manipulation, and multi-agent cooperation scenarios, we demonstrate that NCDPO achieves superior sample efficiency and final performance compared to existing diffusion RL approaches. NCDPO’s ability to handle both continuous and discrete action spaces suggests its potential as a general-purpose policy optimization framework.
Appendix A Self-Imitation Regularizer
Appendix B Additional experimental results
B.1 Original Scores on OpenAI Gym locomotion tasks
Scenario NCDPO (Ours) DPPO AWR IDQL DQL RWR DIPO PPO
Hopper-Medium 3297.3 (47.8) 2566.6 (51.1) 1168.9 (30.5) 2970.2 (5.2) 3200.7 (30.1) 2633.6 (94.0) 2463.1 (127.6) 2609.3
Hopper-Medium-Replay 3345.64 (71.88) 2988.97 (86.90) 2142.30 (183.82) 3076.91 (29.82) 3159.72 (94.85) 2718.27 (79.86) 2834.07 (25.00) 2609.3
Hopper-Medium-Expert 3528.74 (51.89) 2672.64 (135.15) 1062.47 (13.15) 3440.50 (10.79) 3153.66 (733.36) 2964.11 (121.35) 3326.99 (32.73) 2609.3
Walker2d-Medium 4248.8 (137.6) 3632.1 (55.9) 3353.9 (296.9) 3972.8 (17.3) 3405.2 (1322.7) 3238.3 (116.8) 3581.6 (70.3) 3588.5
Walker2d-Medium-Replay 4544.59 (162.98) 3770.52 (154.50) 2719.04 (83.84) 4373.89 (91.66) 3846.26 (1357.97) 2483.04 (215.70) 3180.80 (396.27) 3588.5
Walker2d-Medium-Expert 5060.92 (82.87) 4935.57 (73.28) 4458.68 (195.26) 4863.91 (95.34) 2416.68 (1708.34) 3831.65 (243.76) 4979.96 (50.22) 3588.5
HalfCheetah 7058.8 (635.1) 4758.3 (41.8) 3788.7 (166.5) 4584.4 (45.3) 4459.1 (309.5) 2773.1 (310.0) 4272.4 (33.2) 5783.9
HalfCheetah-Expert 8079.30 (392.21) 4663.23 (92.62) 3723.21 (131.37) 4499.39 (57.61) 4171.97 (465.25) 2218.39 (168.36) 4126.96 (12.14) 5783.9
HalfCheetah-Replay 7000.14 (113.88) 4181.57 (24.59) 3501.08 (40.34) 4295.81 (53.02) 4223.07 (177.28) 1775.98 (101.28) 3362.48 (49.47) 5783.9
Table 3: Mean and standard deviation of original scores across continuous robot control scenarios.
B.2 Experiments with Vision Inputs
We performed evaluation on Square task in robomimic with vision as input. Results demonstrate the effectiveness of NCDPO.
Figure 8: Experimental results for vision inputs.
B.3 Ablation Study
@@
B.4 Further Experiments in OpenAI Gym locomotion tasks.
@@
Regardless of dataset quality, NCDPO consistently outperforms all baselines, as shown in Figures 12 and 13.
@@
B.5 Google Research Football Data Curation
@@ Appendix C Implementation Details and Hyperparameters
Task $\gamma$ $\lambda$ Action Chunk Actor LR Critic LR Actor MLP Size Critic MLP Size Actor MLP Layers $\eta$ Initial Noise Log Std $1/T$ Denoising Step Clone Epochs Clone LR Episode Length Mini-batch Number Environment Max Steps Parallel Environments
@@
Task $\gamma$ $\lambda$ Action Chunk Actor LR Critic LR Actor MLP Size Critic MLP Size Actor MLP Layers Denoising Step Episode Length Mini-batch Size Environment Max Steps Parallel Environments
@@
Table 6: Hyperparameter settings for Baselines in robot control. Experiment is executed using DPPO [20] implementation and hyperparameters. Batch size for all baselines other than DPPO is 1000. For further details, please refer to DPPO paper [20].
Task $\gamma$ $\lambda$ Action Chunk Actor LR Critic LR Actor MLP Size Critic MLP Size Actor MLP Layers $\eta$ Initial Noise Log Std $1/T$ Denoising Step Clone Epochs Clone LR
@@
Task $\gamma$ $\lambda$ Action Chunk Actor LR Critic LR Actor MLP Size Critic MLP Size Actor MLP Layers $\eta$ Initial Log Std $1/T$ Denoising Step
@@
Task $\gamma$ $\lambda$ Action Chunk Actor LR Critic LR Actor MLP Size Critic MLP Size Actor MLP Layers $\eta$ Initial Log Std $1/T$ Denoising Step Clone Epochs Clone LR
@@
Task $\gamma$ $\lambda$ Action Chunk Actor LR Critic LR Actor MLP Size Critic MLP Size Actor MLP Layers $\eta$ Initial Log Std $1/T$ Denoising Step Clone Epochs Clone LR
@@
Task $\gamma$ $\lambda$ Action Chunk Actor LR Critic LR Actor MLP Size Critic MLP Size Actor MLP Layers $\eta$ Initial Log Std $1/T$ Denoising Step Clone Epochs Clone LR

### v4

@@ 6 Experiments
In this section, we provide a comprehensive evaluation of NCDPO across a variety of challenging environments. We first describe the experimental setup in Sec. 6.1, then present results on continuous robot control tasks (Sec. 6.2) and discrete multi-agent coordination tasks (Sec. 6.3). Finally, we report ablation studies in Sec. 6.4 to assess the robustness of NCDPO.
Through extensive experiments aimed to evaluate the capability of NCDPO across diverse continuous control scenarios, our main observations are summarized as follows:
  • NCDPO delivers the best performance: Across all scenarios, NCDPO achieves the best overall performance. Compared with DPPO, it yields an average reward improvement of $30.6\%$ on OpenAI Gym locomotion tasks, $15.4\%$ on Franka Kitchen tasks, and an average success rate improvement of $68.9\%$ on Robomimic tasks with visual inputs (using 4M samples for the transport-vision task).
  • NCDPO achieves the highest sample efficiency. NCDPO consistently attains the highest sample efficiency. Relative to DPPO, it improves sample efficiency by $800.0\%$ in OpenAI Gym locomotion tasks, $34.6\%$ in Franka Kitchen tasks, $97.3\%$ in Robomimic tasks with state inputs, and $313.8\%$ in Robomimic tasks with visual inputs on average. Here, we compare the number of samples required to reach the peak performance achieved by DPPO.
  • NCDPO is robust to the number of denoising steps. NCDPO consistently maintains strong performance and high sample efficiency even when fine-tuning Diffusion Policies with increased numbers of denoising steps.
  • Q-learning methods underperforms in long-horizon, sparse-reward tasks: Q–learning methods underperform and exhibits extreme instability on long-horizon, sparse-reward tasks such as Franka Kitchen and Robomimic, likely due to the increased difficulty of accurately estimating Q-values in these settings. This underscores the importance of on-policy, PPO-style algorithms for tackling such challenging tasks.
For continuous robot control tasks, we compare NCDPO to several baselines: standard Gaussian policies with MLP parameterization; DPPO [24]; DRWR and DAWR [24], which build on reward-weighted regression [21] and advantage-weighted regression [20], respectively; DIPO [34], which uses action gradients as the score function during denoising; and Q-learning–based methods such as IDQL [12] and DQL [32]. To further demonstrate the advantages of our PPO-based approach over Q–learning methods, we also compare against DAC [6] and BDPO [9] on sparse-reward tasks. To ensure a fair comparison, these methods were adapted to an offline-to-online setting by adding a replay buffer.
For the discrete multi-agent environment, Google Research Football, we compare NCDPO with MLP policies trained with Multi-Agent Proximal Policy Optimization (MAPPO) [35]. The MAPPO baseline is initialized via behavior cloning using a cross-entropy loss.
@@ 6.1 Environmental Setup
OpenAI Gym locomotion. We evaluate on standard locomotion benchmarks from OpenAI Gym [3]: Hopper-v2, Walker2D-v2, and HalfCheetah-v2. Pre-trained Diffusion Policies were trained on the D4RL “medium” dataset [7], which contains diverse pre-recorded trajectories. Fine-tuning is performed with dense rewards.
Franka Kitchen. We evaluate on the Franka Kitchen benchmarks [11], where agents must complete four ordered subtasks. Pretraining datasets are complete, mixed, and partial from D4RL [7]. Fine-tuning uses sparse rewards (reward = 1 when a subtask is completed).
Robomimic. We evaluate on Robomimic manipulation tasks [18]: Lift, Can, Square, and Transport. To ensure temporal consistency, we apply action chunking with chunk size 4 for Lift, Can, and Square, and chunk size 8 for Transport, following [24]. All Robomimic tasks are fine-tuned with sparse success/failure rewards.
Google Research Football. To evaluate performance in large discrete joint-action spaces, we test three multi-agent scenarios: 3 vs 1 with Keeper, Counterattack Hard, and Corner. We adopt a centralized control strategy: a single Diffusion Policy generates joint actions for all agents simultaneously. Base Diffusion Policies are pre-trained to output one-hot vectors representing ground-truth actions. Because no public dataset exists for these scenarios, we construct a pretraining dataset by training multiple MAPPO agents [35] with different random seeds and early-stopping schedules; this produces a diverse set of trajectories with varying success rates and tactics.
@@ 6.2 Evaluation on Continuous Robot Control Tasks
State Inputs. Experimental results in Fig. 3 and Table 1 demonstrate that NCDPO consistently achieves the strongest performance, highest robustness, and superior sample efficiency across all evaluated tasks with state input. These findings highlight substantial improvements over DPPO and other baselines.
Notably, the Transport scenario is exceptionally challenging: it requires coordinating two robot arms across multiple complex sub-tasks, with extremely long horizons and sparse rewards. To the best of our knowledge, no reinforcement learning algorithm other than DPPO—whether off-policy or on-policy—has achieved a success rate above 50% on this task. The strong performance of NCDPO in this setting highlights its robustness and broad applicability.
Figure 3: Performance comparisons on OpenAI Gym locomotion, Franka Kitchen, and Robomimic tasks. Results are averaged over three seeds. NCDPO (ours) achieves the strongest performance.
Vision Inputs. We also evaluate with vision inputs on the two most challenging Robomimic tasks (Square and Transport). Results are demonstrated in Fig. 4. Remarkably, NCDPO achieves substantial performance gains and attains roughly three times higher sample efficiency than DPPO on these visual tasks, underscoring its surprising effectiveness and strong potential for real-world applications.
Figure 4: Performance comparison on Robomimic visual tasks. Results are averaged over three seeds. NCDPO (ours) achieves the strongest performance. VIT-MLP uses a Vision Transformer encoder and a Gaussian policy parameterize by MLP.
@@ 6.3 Evaluation on Discrete Multi-agent Coordination Tasks
We next evaluate NCDPO on Google Research Football, a cooperative multi-agent benchmark. As shown in Fig. 5 and Table 2, NCDPO consistently outperforms the MAPPO (MLP) baseline across all three scenarios. These results highlight the strength of Diffusion Policies in modeling complex and diverse demonstration data, as well as the effectiveness of the DP+NCDPO fine-tuning procedure during fine-tuning phase. Moreover, they demonstrate the versatility of NCDPO, showing its effectiveness not only lies in continuous robot control tasks but also in environments with large discrete action spaces.
@@
6.4 NCDPO Is Robust to the Number of Denoising Steps
We study the effect of varying the number of denoising steps in the diffusion model. Results in Fig. 6 indicate that NCDPO is robust to this choice across tasks. We hypothesize that this robustness stems from how gradients propagate through time in the diffusion process, which yields more accurate gradient estimates during fine-tuning.
Figure 6: Ablation study on denoising steps in OpenAI Gym locomotion tasks and Robomimic tasks.
7 Conclusion and Limitations
We present NCDPO, a novel approach for fine-tuning Diffusion Policies through Proximal Policy Optimization that exhibits strong performance across continuous and discrete control domains. Our key innovation lies in backpropagating gradients through diffusion timesteps by reformulating the diffusion denoising process as a noise-conditioned stochastic policy. Through extensive experiments across locomotion, manipulation, and multi-agent cooperation scenarios, we demonstrate that NCDPO achieves superior sample efficiency and final performance compared to existing diffusion RL approaches. NCDPO’s ability to handle both continuous and discrete action spaces suggests its potential as a general-purpose policy optimization framework.
Appendix A Self-Imitation Regularizer
Appendix B Additional experimental results
B.1 Scores for Continuous Robot Control Tasks
Scenario NCDPO (Ours) DPPO AWR IDQL DQL RWR DIPO DAC BDPO QSM MLP
hopper 3297.3 (47.8) 2566.6 (51.1) 1168.9 (30.5) 2970.2 (5.2) 3200.7 (30.1) 2633.6 (94.0) 2463.1 (127.6)
hopper-replay 3345.6 (71.9) 2989.0 (86.9) 2142.3 (183.8) 3076.9 (29.8) 3159.7 (94.8) 2718.3 (79.9) 2834.1 (25.0)
hopper-expert 3528.7 (51.9) 2672.6 (135.1) 1062.5 (13.1) 3440.5 (10.8) 3153.7 (733.4) 2964.1 (121.3) 3327.0 (32.7)
walker2d 4248.8 (137.6) 3632.1 (55.9) 3353.9 (296.9) 3972.8 (17.3) 3405.2 (1322.7) 3238.3 (116.8) 3581.6 (70.3)
walker2d-replay 4544.6 (163.0) 3770.5 (154.5) 2719.0 (83.8) 4373.9 (91.7) 3846.3 (1358.0) 2483.0 (215.7) 3180.8 (396.3)
walker2d-expert 5060.9 (82.9) 4935.6 (73.3) 4458.7 (195.3) 4863.9 (95.3) 2416.7 (1708.3) 3831.7 (243.8) 4980.0 (50.2)
halfcheetah 7058.8 (635.2) 4758.3 (41.8) 3788.7 (166.5) 4584.4 (45.3) 4459.1 (309.5) 2773.1 (310.0) 4272.4 (33.2)
halfcheetah-replay 7000.1 (113.9) 4181.6 (24.6) 3501.1 (40.3) 4295.8 (53.0) 4223.1 (177.3) 1776.0 (101.3) 3362.5 (49.5)
halfcheetah-expert 8079.3 (392.2) 4663.2 (92.6) 3723.2 (131.4) 4499.4 (57.6) 4172.0 (465.3) 2218.4 (168.4) 4127.0 (12.1)
kitchen-complete-v0 4.0 (0.0) 3.8 (0.1) 1.6 (0.4) 0.0 (0.0) 3.4 (0.3)
kitchen-mixed-v0 4.0 (0.0) 3.3 (0.6) 1.7 (0.7) 2.2 (0.2) 1.8 (0.1)
kitchen-partial-v0 3.6 (0.5) 3.0 (0.0) 1.1 (1.0) 2.5 (1.1) 1.9 (0.1)
lift 100.0 (0.0) 99.7 (0.3) 93.3 (1.8) 99.2 (1.0) 99.8 (0.3) 97.5 (0.5) 97.3 (0.8) 97.2 (3.3) 100.0 (0.0) 94.7 (9.2) 92.5 (0.0)
can 99.3 (1.2) 99.0 (1.0) 33.8 (3.2) 94.5 (3.1) 0.3 (0.6) 90.7 (0.8) 52.8 (5.1) 12.0 (2.8) 0.0 (0.0) 94.7 (9.2) 84.8 (2.5)
square 87.3 (4.5) 87.0 (2.3) 40.3 (8.5) 80.0 (5.0) 0.0 (0.0) 74.8 (2.6) 25.3 (4.5) 52.2 (3.9) 0.0 (0.0) 0.7 (0.6) 64.5 (9.5)
transport 96.7 (2.3) 91.3 (2.9) 11.2 (3.5) 0.5 (0.9) 0.0 (0.0) 0.0 (0.0) 0.2 (0.3) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0) 0.0 (0.0)
square-vision 85.5 (2.1) 54.8 (3.5) 9.7 (1.6)
transport-vision (4M) 91.3 (4.6) 50.2 (3.3) 0.8 (0.6)
Table 1: Mean and standard deviation of performance across robot control tasks.
@@ B.1 Scores for Continuous Robot Control Tasks
B.2 Ablation Study
@@
B.3 Further Experiments in OpenAI Gym locomotion tasks.
@@
Regardless of dataset quality, NCDPO consistently outperforms all baselines, as shown in Figures 10 and 11.
@@
B.4 Google Research Football Data Curation
@@ Appendix C Implementation Details and Hyperparameters
Task $\gamma$ $\lambda$
Action
Chunk
Actor
LR
Critic
LR
Actor
MLP Size
Critic
MLP Size
Actor
MLP Layers
$\eta$
Initial Noise
Log Std
$1/T$
Denoising
Step
Clone
Epochs
Clone
LR
Episode
Length
Mini-batch
Number
Environment
Max Steps
Parallel
Environments
@@
Task $\gamma$ $\lambda$
Action
Chunk
Actor
LR
Critic
LR
Actor
MLP Size
Critic
MLP Size
Actor
MLP Layers
Denoising
Step
Episode
Length
Mini-batch
Size
Environment
Max Steps
Parallel
Environments
@@
Table 5: Hyperparameter settings for Baselines other than BDPO and DAC in robot control tasks. Experiment is executed using DPPO [24] implementation and hyperparameters. Batch size for all baselines other than DPPO is 1000. For further details, please refer to DPPO paper [24].
Task
Rollout Samples
Rollout Interval
Warm Up
Envs
Critic Rho
Critic Samples
Replay
Buffer Size
Kitchen-DAC 10 5000 - 50 0 10 1e6
Kitchen-BDPO 10 5000 50000 50 0 10 1e6
Robomimic-DAC 1 2000 - - - 10 1e6
Robomimic-BDPO 1 2000 20000 50 0 10 1e6
Table 6: Hyperparameters for DAC and BDPO on Kitchen and Robomimic tasks. Other hyperparameters and implementation is from Gao & Cao [8]. Here we use 1 sample for rollout because we find it difficult to learn an accurate Q-estimation, which leads to severely degraded policy quality when number of samples is greater than 1.
Task $\gamma$ $\lambda$
Action
Chunk
Actor
LR
Critic
LR
Actor
MLP Size
Critic
MLP Size
Actor
MLP Layers
$\eta$
Initial Noise
Log Std
$1/T$
Denoising
Step
Clone
Epochs
Clone
LR
@@
Task $\gamma$ $\lambda$
Action
Chunk
Actor
LR
Critic
LR
Actor
MLP Size
Critic
MLP Size
Actor
MLP Layers
$\eta$
Initial
Log Std
$1/T$
Denoising
Step
@@
Task $\gamma$ $\lambda$
Action
Chunk
Actor
LR
Critic
LR
Actor
MLP Size
Critic
MLP Size
Actor
MLP Layers
$\eta$
Initial
Log Std
$1/T$
Denoising
Step
Clone
Epochs
Clone
LR
@@
Task $\gamma$ $\lambda$
Action
Chunk
Actor
LR
Critic
LR
Actor
MLP Size
Critic
MLP Size
Actor
MLP Layers
$\eta$
Initial
Log Std
$1/T$
Denoising
Step
Clone
Epochs
Clone
LR
@@
Task $\gamma$ $\lambda$
Action
Chunk
Actor
LR
Critic
LR
Actor
MLP Size
Critic
MLP Size
Actor
MLP Layers
$\eta$
Initial
Log Std
$1/T$
Denoising
Step
Clone
Epochs
Clone
LR